# College of Emergency Preparedness, Homeland Security and Cybersecurity

### v1

@@ 5 Conclusion
In this study, we applied network analysis to Attention-Deficit/Hyperactivity Disorder (ADHD) by constructing a multimodal knowledge graph that integrates data from diverse sources and expert knowledge. This structured representation maps intricate relationships between various factors associated with ADHD and offers deeper insights into its etiology and manifestation. Dr. Faraone’s extensive research was instrumental in informing our approach, providing validated measures that enhanced the relevance and robustness of our study. By building upon his contributions, we grounded our analysis in the most authoritative understanding of ADHD, further strengthening the credibility of our findings.
Our methodology leverages advanced techniques like contextual proximity calculations and node embeddings using Large Language Models (LLMs) to extract and structure ADHD-related knowledge. This approach reduced redundancy, captured subtle semantic connections, and employed clustering techniques to merge semantically similar concepts, resulting in a streamlined and nuanced knowledge graph. Despite using a relatively small dataset, the graph uncovered an extensive network of ADHD-related concepts, revealing its broad connections across various topics and disciplines. The graph’s potential for expansion further underscores its value as a research tool.
We also identified areas for improvement, such as eliminating irrelevant nodes to enhance interpretability and reduce computational complexity. This refinement process highlighted challenges in data extraction and preprocessing, suggesting areas for methodological improvements to optimize the graph further. The knowledge graph maps ADHD’s interdisciplinary connections, revealing intersections with conditions like depression, schizophrenia, and unexpected areas like lung cancer. Additionally, the absence of treatment-related K-cores points to gaps in the dataset, suggesting that therapeutic strategies may be underrepresented, requiring more refined techniques to distinguish core ADHD concepts.
@@
Appendix
Listing 1: System prompt that used for node and edge extraction
You are a network graph maker who extracts terms and their relations from a given context.
You are provided with a context chunk (delimited by ‘‘‘)
Your task is to extract the ontology of terms mentioned in the given context.
These terms should represent the key concepts as per the context.
Thought 1: While traversing through each sentence, Think about the key terms mentioned in it. Terms must only include entity, location, organization, person, condition, documents, service, concept, date. Terms should be as atomistic as possible.
Thought 2: Think about how these terms can have one on one relation with other terms. Terms that are mentioned in the same sentence or the same paragraph are typically related to each other. Terms can be related to many other terms.
Thought 3: Find out the relation between each such related pair of terms.
[…]
atomoxetine & entity & fda & organization & is approved by \\
[…]
atomoxetine & entity & adhd & condition & treats \\
[…]
atomoxetine & entity & children & person & is approved for the treatment of in \\
[…]
atomoxetine & entity & adolescents & entity & is approved for the treatment of in \\
[…]
atomoxetine & entity & adults & person & is approved for the treatment of in \\
[…]
atomoxetine & entity & norepinephrine & entity & works by increasing the amount of certain available for cells to communicate with each other in \\
[…]
atomoxetine & entity & dopamine & entity & works by increasing the amount of certain available for cells to communicate with each other in \\
[…]
atomoxetine & entity & stimulants & concept & unlike, is not a controlled substance \\
[…]
atomoxetine & entity & several-month supply of medication & other & makes it easier to obtain a \\
[…]

### v2

@@ 5 Conclusion
In this study, we applied network analysis to Attention-Deficit/Hyperactivity Disorder (ADHD) by constructing a multimodal KG that integrates data from diverse sources and expert knowledge. This structured representation maps intricate relationships between various factors associated with ADHD and offers deeper insights into its etiology and manifestation. Dr. Faraone’s extensive research was instrumental in informing our approach, providing validated measures that enhanced the relevance and robustness of our study. By building upon his contributions, we grounded our analysis in the most authoritative understanding of ADHD, further strengthening the credibility of our findings.
Our methodology leverages advanced techniques like contextual proximity calculations and node embeddings using Large Language Models (LLMs) to extract and structure ADHD-related knowledge. This approach reduced redundancy, captured subtle semantic connections, and employed clustering techniques to merge semantically similar concepts, resulting in a streamlined and nuanced KG. Despite using a relatively small dataset, the graph uncovered an extensive network of ADHD-related concepts, revealing its broad connections across various topics and disciplines. The graph’s potential for expansion further underscores its value as a research tool.
We also identified areas for improvement, such as eliminating irrelevant nodes to enhance interpretability and reduce computational complexity. This refinement process highlighted challenges in data extraction and preprocessing, suggesting areas for methodological improvements to optimize the graph further. The KG maps ADHD’s interdisciplinary connections, revealing intersections with conditions like depression, schizophrenia, and unexpected areas like lung cancer. Additionally, the absence of treatment-related K-cores points to gaps in the dataset, suggesting that therapeutic strategies may be underrepresented, requiring more refined techniques to distinguish core ADHD concepts.